Gina rented a moon-walk for 5 hours at a total cost of $235. Lori rented the same moon-walk for 3 hours at a total cost of $165. The cost of renting is based upon a flag set up fee plus a rental rate per hour. How much is the set up fee? What is the hourly rental fee?
Answer by ankor@dixie-net.com(22740)

let f = the flat set-up fee
let r = the price per hr of use
:
Write an equation for each statement
:
Gina rented a moon-walk for 5 hours at a total cost of $235.
5r + f = 235
Lori rented the same moon-walk for 3 hours at a total cost of $165.
3r + f = 165
We can use elimination here
5r + f = 235
3r + f = 165
----------------Subtraction eliminates f, find r
2r + 0 = 70
r = 70/2
r = $35 an hr rental
then find f
3(35) + f = 165
105 + f = 165
f = 165 - 105
f = $60 is the set-up fee
:
:
Check this in the Gina's equation
5(35) + 60 =
175 + 60 = $235


How much is the set up fee? What is the hourly rental fee?